# Determination of impact parameter for CEE with digi-input neural networks

### v3

@@ Abstract
The impact parameter characterizes the centrality in nucleus-nucleus collision geometry. The determination of impact parameters in real experiments is usually based on the reconstructed tracks and hits, or the derived event-level observables. For the scheduled Cooler-storage-ring External-target Experiment (CEE), the low beam energy results in a reduction of dependency between the impact parameter and charged particle multiplicity, which decreases the validity of the explicit determination methods. This work implements a few neural networks that directly take the digitized signals from the external Time-of-flight detectors as input. The model with the best performance has shown a mean absolute error of 0.479 fm with the simulated U-U collisions at 0.5 AGeV. Performances of these models with digi inputs are compared with reference models with phase space inputs, showing the capacity of neural networks in handling the original but potentially interrelated digitized signal information.
@@ 1 Introduction
Relativistic nucleus-nucleus collision is one of the main approaches to studying the Equation of State (EoS) of the nuclear matter[1]. By designing the collision system and energy, the compressed matter is expected to reach a specific state in the Quantum Chromodynamics (QCD) phase diagram[2]. Although such a state evolution lasts only for tens of fm/$c$, results and evidence on physics issues such as critical point of phase transition[3], Quark-Gluon Plasma[4], color-superconductivity[5] can be made through the detection and analysis of the final-state particles. There are many colliders[6, 7, 8] or fixed-target facilities[9, 10] around the world whose detectors are built for the acceptance of collision products such as the charged particles, neutrons, gamma rays, and projectile fragments. The Cooler-storage-ring External-target Experiment (CEE)[11] is a fixed-target heavy-ion experiment that will run in an energy range of 0.5-1.2 AGeV. Time-of-flight (TOF) technique[12], which requires cooperation of tracking and timing detectors, is practiced for charged particle identifications. Figure 1 shows the detector layout of the CEE. Since there is a large variation of particle momentum with polar angles, two arrays of detectors are designed for identifying the particles at the front ($\theta$<25°) and large angle ($\theta$>25°). The Time Projection Chamber (TPC)[13] and the inner TOF (iTOF)[14] system are installed inside the dipole magnet and cover the intermediate rapidity zone. As front-angle detectors, the Multi-Wire Drift Chamber (MWDC)[15] and the external TOF (eTOF)[16, 17] wall are placed downstream of the magnet for a longer distance. Figure 2 shows the phase space distribution, i.e. the particle distribution presented by the reduced rapidity $y_{0}$ and the reduced transverse momentum $p_{\rm T}/m$, where the acceptances of the two detector arrays can be recognized. A T0 detector[18] is also included for start time measurements. A Zero-Degree Calorimeter (ZDC)[19] is placed a the most downstream, measuring the baryons for the reconstruction of event plane and centrality.
@@
Figure 2: Phase space distribution of CEE. The size of the pixels stands for the particle density, and the contour shows the coverage of the forward and transverse detectors.
@@
Given a fixed beam energy, the centrality serves as the most descriptive geometric quantity which determines the initial state of a collision. An equivelent description of centrality is the impact parameter $b$, which is defined as the perpendicular distance between the centers of both nuclei along the trajection. More central collisions result in more interacting nucleons known as participants, and compressed nuclear matter with higher density. There are several EoS-related research topics, such as the collective and elliptical flow, and symmetry energy, which need the recognition of the central, semi-peripheral, and peripheral collisions[20, 21]. The most common method in experiments is to measure the distribution of a single observable with a strong correlation to the impact parameter. Then the distribution can be processed by a sharp cut-off or Bayesian methods which give discrimination strategies[22, 23, 24, 25]. Such observables include the total charged multiplicity, the transverse momentum, the number of intermediate-mass fragments, the total number of participant neutrons, etc. The fruitful collection of the above features also inspired the multivariate analysis, which turned the centrality determination into a typical instance of feature recognition. Since the 1990s, there have been works that apply Support Vector Machine (SVM)[26], Gradient Boosting Decision Tree (GBDT)[27], and several kinds of neural networks[28, 29, 30] to the regression of the impact parameters. Such works have two main characteristics: the regression or classification model should be trained with simulated data which can provide the target ‘TRUE’ impact parameter; the input variables to the models are the highly correlated observables obtained through event reconstruction.
@@
The development of Artificial Intelligence and Deep Neural Networks (DNN) has expanded the capabilities of regression models in non-linear tasks and made it possible for models to handle data with primitive features. For recent works on centrality determination, tracks and hits of the final state particles, extracted from the detectors, have been considered as direct features. F. Li, et al translated the particle production into a phase space image in a 30×30 shape and implemented the Convolution Neural Network (CNN)[31] to regress the impact parameters[32]. Then they tested the precision of the method under the Au-Au collision data in an intermediate low beam energy range of 0.2-1.0 AGeV. M. O. Kuttan, et al were dedicated to developing a CNN-based model called PointNet which could handle the variable length and commutable input of hits and tracks[30]. The whole structure was applied to the regression of the impact parameter in Au-Au collisions at 10 AGeV.
@@
The above works inspire the impact parameter determination of CEE. We propose the regression with DNNs using the channel response of the eTOF detectors as input. It is shown from the tests that the utilization of the detector data results in even higher prediction powers on the impact parameter than the common event-level features, such as phase space distribution. This paper is organized as follows: Section 2 describes the task of impact parameter determination in the context of CEE, basic information about the eTOF wall, and how the features from the eTOF detectors are extracted. Section 3 describes how the simulation is designed and carried out, including the data preparation and the DNN models. Section 4 shows the simulation results of the models. A summary is made in the last section.
@@ 2 Description of the task
The typical collision system for CEE is U-U at 0.5 AGeV. The collision energy is below the required threshold for most mesons and resonances, leaving the majority of the charged products as baryons like protons, deuterons, tritons and heavier ions[17]. As a result, the event fluctuation for central collisions is high, which deteriorates the centrality dependence with multiplicity. From the example of Au-Au collisions[30], we can read the precision from polynomial fit around 1 fm for events at 10 AGeV with $b=1$ fm. However, such a prediction is not valid for events at 1 AGeV because the prediction stops around 2 fm. Given the expectation of high prediction uncertainty in the CEE condition, many strategies for centrality determination have been considered, for example, by using the angle distribution of energy deposition in ZDC, or by using reconstructed tracks of the full system. It is expected that strategies based on different detectors as data sources can be combined into a powerful final predictor.
[…]
@@
Because of the low collision energy and the limited front-angle coverage, the monotonicity between the accepted particle multiplicity and impact parameter loses totally for the eTOF wall, as shown in Figure 4. The multiplicity starts to drop when the impact parameter goes under 5 fm, likely due to the increase of average transverse momentum in central collisions. Therefore, it is important to consider multiple features when using eTOF as the data source. In other words, not only the accepted hit muliplicity, but also their temporal and spatial features should be included. Such features can be easily extracted from the digitized signals, i.e. the recorded leading and trailing time mentioned above. The average time of both ends of the strip represents the hit time, and the differential time represents the hit position along the strip. The channel layout of the eTOF wall can be considered as a gridding of the phase space. Therefore, it is reasonable to believe that the digi-input strategy will not lose either the information or the granularity in the context of centrality determination.
Figure 4: Multiplicity distribution of eTOF with respect to impact parameters.
This work gives a simulation study using the simulated collisions and the resulting digitized signals from eTOF. However, it is important to maintain the validity of the methods and models when using the experimental signals in future practices. At least two factors should be concerned on the quality of the experimental data. The first is the correct event assembly, which requires that the acquired signals are from a single collision. In CEE condition, over 98% of the in-event charged particles reach the TOF detectors within a time window of 25 ns. Accordingly, the time windows for triggering and data acquisition are 75 ns and 1 us respectively. Given the interaction rate of 10 kHz, the channel response of eTOF in the same event can be organized correctly. The second factor is the stability of the data distribution, such as the reconstructed time and position. It will introduce prediction errors to the model if the experimental data distribution is biased or unstatic. During the operation of CEE, parameters from both the detectors (signal propagation speed, time resolution, etc.) and the electonics (processing time of readout, triggering, etc.) are stable with time. Given a feature distribution independent with instrument conditions, proper normalization can be carried out to mininize the divergence between the simulation and experimental data distributions.
@@ 3.1 Dataset generation
Datasets in this paper are generated following the procedures below: Firstly, collision data is generated which gives the produced steady particles and their initial kinetic attributes. The physics model of the collision is Isospin-dependent Quantum Molecular Dynamics (IQMD)[35], and the coalescence of nucleons is processed after the IQMD calculation since ion yield is significant in the CEE energy. Then, CeeRoot[36], a platform based on the FairRoot[37] framework, carries out the simulation which organizes the collision data together with detector geometry and the Geant4 transport model. Finally, with proper pre-processing, we obtain the dataset for model training and testing.
[…]
@@
where $t_{\rm hit}$, $x_{\rm hit}$ are the hit time and position, $x_{\rm end}$ is the position of the strip end, $v_{\rm prop}$ is the signal propagation speed along the strip. $\Delta t_{\rm ele}$ is the time delay of the electronics, which is sampled from uniform distribution [0, 5) ns and constant for each specific channel. $\Delta t_{\rm reso}$ is sampled from Gaussian distribution with 100 ps standard deviation, representing the time resolution. It is inferred from the data normalization that the time delay effect can be eliminated in the model regression and prediction, which is very helpful because of the complexity of the measurement of channel time delay. As the time-slewing effect has not been implemented in the digitization method, the 100 ps smearing is chosen according to the measurement before the time-slewing correction in the cosmic tests[38].
@@
We simulated 1.21 million events of U-U collisions at 0.5 AGeV. The events are distributed uniformly from impact parameter b=0.0 fm to b=12.0 fm with 0.1 fm step. The total events are randomly split into the train and test set with an 8:2 ratio. Two datasets are generated based on the test set: one is the so-called $b{\rm d}b$ test set with 120k events, and the other is the uniform test set with all the 242k events. The $b{\rm d}b$ test set is generated following the collision geometry where the differential cross section is proportional to the impact parameter: ${\rm d}\sigma\propto b{\rm d}b$. It is used for evaluating the overall performance of the models because it is close to the realistic collisions; the uniform test set is for the assessment of performance in the function of the impact parameters. The train set has 968k events in which 80% are used for model fitting and 20% are used for early-stop decision (see Section 3.3) after each training epoch.
@@
For each simulated event, we collect the features from the total 672 readout strips, including the hit flag, hit time, and hit position. The hit flag is set to 1 if there are signals generated from both ends of the strip. The hit time is calculated as the average time of both ends and the hit position is the half difference of the both-end time values. Such calculation, despite its simplicity, provides features with clearer physics meaning. For strips with no hit detected, the hit flag is set to zero and the time and position are set to the local average value from the training dataset. The choice of data imputation here is made in order to obtain zero values after the normalization of the time and position features. The purpose of keeping zero values for the non-fired nodes is to implicitly stop the forward propagation of their features. The first dense layer is designed to exclude the intercept term, so that the non-fired nodes with zero input values always result in zero outputs. The absence of trainable intercept have no effect on the capability of the dense layer because the weight on the hit flag plays an equivalent role. Moreover, the zero-value nodes have no function in the attention layer when the linear combination of the nodes are calculated with their attributes. Figure 5 shows the counting rate distribution of the 672 readout strips for the eTOF wall. A descending conting rate is shown due to the collision in fix-target mode.
@@
Figure 5: Counting rate distribution of 672 readout strips for eTOF wall. The color bar is in Hz.
@@
To compare the model performance with other work, we generate the phase space dataset with 30x30 binning. The range of the phase space used for binning is defined as $y_{0}\in$[-0.5,1.5], $p_{\rm T}/m\in$[0,1.5]. Figure 6 shows the cumulative images of the phase space with impact parameters, where the evolution of the image shape can be easily recognized. However, for a single event, features get more indistinct due to the event fluctuation, as demonstrated in Figure 6. Two sets of phase space data, with acceptances covered by CEE and only eTOF, have been prepared for a comparison study.
@@
Figure 6: (a) The resampling phase space images in the acceptance of CEE with different impact parameters. Each image is filled with 10k events with the impact parameter shown in the legend. The binning of the image is 30×30. (b) Image of an event with $b=6$ fm.

### v4

@@ Abstract
The impact parameter characterizes the centrality in nucleus-nucleus collision geometry. The determination of impact parameters in real experiments is usually based on the reconstructed particle attributes or the derived event-level observables. For the scheduled Cooler-storage-ring External-target Experiment (CEE), the low beam energy reduces correlation between the impact parameter and charged particle multiplicity, which decreases the validity of the explicit determination methods. This work investigates a few neural network-based models that directly take the digitized signals from the external Time-of-flight detectors as input. The model with the best performance shows a mean absolute error of 0.479 fm with simulated U-U collisions at 0.5 AGeV. The performances of the models implemented with digi inputs are compared with reference models with phase space inputs, showing the capability of neural networks to handle the original but potentially interrelated digitized signal information.
@@ 1 Introduction
Relativistic nucleus-nucleus collision is one of the main approaches to studying the Equation of State (EoS) of the nuclear matter[1]. By designing the collision system and energy, the compressed matter is expected to reach a specific state in the Quantum Chromodynamics (QCD) phase diagram[2]. Although such a state evolution lasts only for tens of fm/$c$, results and evidence on physics issues such as critical point of phase transition[3], Quark-Gluon Plasma[4], color-superconductivity[5] can be made through the detection and analysis of the final-state particles. There are many colliders[6, 7, 8] or fixed-target facilities[9, 10] around the world whose detectors are built for the acceptance of collision products such as charged particles, neutrons, gamma rays, and projectile fragments. The Cooler-storage-ring External-target Experiment (CEE)[11] is a fixed-target heavy-ion experiment running in an energy range of 0.5-1.2 AGeV. The time-of-flight (TOF) technique[12], which requires the cooperation of tracking and timing detectors, is practiced for charged particle identification. Figure 1 shows the detector layout of the CEE. Since there is a considerable variation of particle momentum with polar angle, two arrays of detectors are designed for identifying the particles at the front ($\theta$<25°) and at large angles ($\theta$>25°). The Time Projection Chamber (TPC)[13] and the inner TOF (iTOF)[14] systems are installed inside the dipole magnet and cover the intermediate rapidity zone. As front-angle detectors, the Multi-Wire Drift Chamber (MWDC)[15] and the external TOF (eTOF)[16, 17] wall are placed downstream of the magnet, allowing longer flight distance for better resolution in particle identification. Figure 2 shows the phase space distribution, i.e., the particle distribution as a function of the reduced rapidity $y_{0}$ and the reduced transverse momentum $p_{\rm T}/m$, where the acceptances of the two detector arrays can be recognized. A T0 detector[18] is also included for start time measurements. A Zero-Degree Calorimeter (ZDC)[19] is placed behind the eTOF to measure the baryons used for the reconstruction of event plane and centrality.
@@
Figure 2: Phase space distribution of CEE. The size of the pixels stands for the particle density, and the contours show the coverage of the forward and transverse detectors.
@@
Given a fixed beam energy, centrality is the most descriptive geometric quantity determining a collision’s initial state. An equivalent description of centrality is the impact parameter $b$, defined as the perpendicular distance between the trajectories of both nuclei before their collision. Central collisions result in a large number of interacting nucleons, known as participants, and compressed nuclear matter with high density. There are several EoS-related research topics, such as the collective and elliptical flow and symmetry energy, which need the recognition of the central, semi-peripheral, and peripheral collisions[20, 21]. The most common method in experiments is to measure the distribution of a single observable with a strong correlation with the impact parameter. Then, the distribution can be processed by a sharp cut-off or Bayesian methods, which give discrimination strategies[22, 23, 24, 25]. Such observables include the total charged multiplicity, the transverse momentum, the number of intermediate-mass fragments, the total number of participant neutrons, etc. The fruitful collection of the above features also inspired the multivariate analyses, which turned the centrality determination into a typical instance of feature recognition. Since the 1990s, there have been works that apply Support Vector Machine (SVM)[26], Gradient Boosting Decision Tree (GBDT)[27], and several kinds of neural networks[28, 29, 30] to the regression of the impact parameter. Such works have two main characteristics: the regression or classification model is trained with simulated data, which can provide the target ‘TRUE’ impact parameter, and the input variables to the models are the highly correlated observables obtained through event reconstruction.
@@
The development of Artificial Intelligence and Deep Neural Networks (DNN) has expanded the capabilities of regression models in non-linear tasks and made it possible for models to handle data with primitive features. In recent works on centrality determination, tracks and hits of the final state particles extracted from the detectors have been considered direct features. F. Li et al. translated the particle production into a phase space image in a 30$\times$30 shape of pixels and implemented a Convolution Neural Network (CNN)[31] based model to regress the impact parameter[32]. Then, they tested the precision under the Au-Au collision data in an intermediate low beam energy range of 0.2-1.0 AGeV. M. O. Kuttan et al. were dedicated to developing a CNN-based model called PointNet, which could handle the variable length and commutable input of hits and tracks[30]. The whole structure was applied to the regression of the impact parameter in Au-Au collisions at 10 AGeV.
@@
The above works inspired the impact parameter determination of CEE. We propose the regression with DNNs using the channel response of the eTOF detectors as input. Tests showed that using data from the detector leads to more accurate predictions of the impact parameter compared to using common event-level features like phase space distribution. This paper is organized as follows: Section 2 describes the task of impact parameter determination in the context of CEE, provides basic information about the eTOF wall, and accounts for how the features from the eTOF detectors are extracted. Section 3 describes how the simulation is designed and carried out, including the data preparation and the DNN models. Section 4 is dedicated to the discussion of results. A summary is made in the last section.
@@ 2 Description of the task
The typical collision system for CEE is U-U at 0.5 AGeV. The collision energy is below the required threshold for most mesons and resonances, leaving most charged products as baryons like protons, deuterons, tritons, and heavier ions[17]. As a result, event-by-event fluctuations for central collisions are high, which limit the centrality dependence on multiplicity[30]. Given the expectation of high prediction uncertainty in the CEE condition, many strategies for centrality determination have been considered, such as using the angle distribution of energy deposition in ZDC or reconstructed tracks of the entire system. It is expected that strategies based on different detectors as data sources can be combined into a decisive final predictor.
The eTOF wall is one of the best detector candidates for centrality determination thanks to the extensive coverage (3.2$\times$1.6 m2), high granularity (1344 channels in total), and fast signal shaping. The eTOF wall comprises 24 Multigap Resistive Plate Chambers (MRPC) with 10 gas gaps of 0.25 mm thickness. Figure 3 shows the detector layout of the eTOF wall, where the MRPCs cover an active area of 3.2$\times$1.6 m2. The eTOF wall contains 672 horizontal strips, read out from both ends. They are in the same size of 48$\times$1.5 cm2 and arranged in an interval of 1.7 cm in every detector. The 6 smaller MRPCs in the inner region of the eTOF wall have 16 readout strips, while the others have 32. The eTOF wall is divided into 7 columns, which results in a 96$\times$7 layout of readout strips. The primitive signals, induced onto the readout strips by the avalanches in MRPC’s gas gaps, are amplified and discriminated by the NINO-based[33] Front-End Electronics (FEE)[34]. The output signals are recorded by the Time-to-Digital Module[34] for their leading and trailing time when the amplified signals cross a threshold (i.e., 150 mV for eTOF FEEs). The expected efficiency and time resolution of the eTOF MRPCs are better than 95% and 60 ps, respectively, and have been verified on real-size prototypes in the cosmic and beam tests.
@@
Due to the low collision energy and limited coverage at frontal angles ($\theta$<25°), there is a complete loss of monotonic relationship between accepted particle multiplicity and impact parameter for the eTOF wall, as illustrated in Figure 4. The observed positive correlation between multiplicity and impact parameter, specifically for impact parameters less than 5 fm, is probably due to the particle acceptance influenced by transverse momentum in central collisions. Conversely, the negative correlation for $b$ exceeding 5 fm is resulted from the diminishing number of participants. Therefore, it is essential to consider multiple features when using eTOF as a data source. In other words, not only the accepted hit multiplicity but also their temporal and spatial features should be included. Such features can be easily extracted from the digitized signals, i.e., the recorded leading and trailing-edge times mentioned above. The average time of both ends of the strip represents the hit time, and the differential time represents the hit position along the strip. The channel layout of the eTOF wall represents a gridding of the phase space. Therefore, it is reasonable to believe that the digi-input strategy will not lose the information or the granularity in the context of centrality determination.
Figure 4: Multiplicity distribution of eTOF as a function of impact parameter.
This work presents a procedure which takes digitized signals from eTOF and gives event-by-event predictions of impact parameters. However, it is important to maintain the validity of the methods and models when using the experimental signals in future practices. At least two factors should be considered on the quality of the experimental data. The first is the correct event assembly, which requires that the acquired signals are sourced from a single collision. In CEE condition, over 98% of the in-event charged particles reach the TOF detectors within a time window of 25 ns. Accordingly, the time windows for triggering and data acquisition are 75 ns and 1 us, respectively. Given the interaction rate of 10 kHz, the channel response of eTOF in the same event can be organized correctly. The second factor is the stability of the data distribution, such as the reconstructed time and position. Should the experimental data distribution be biased or unstatic, model prediction errors would result. During the operation of CEE, parameters from both the detectors (signal propagation speed, time resolution, etc.) and the electronics (processing time of readout, triggering, etc.) are stable with time. Given a feature distribution independent from instrument conditions, proper normalization can be carried out to minimize the differences between the simulation and experimental data distributions. The simulation study utilizes ground truth information to evaluate model performances. In practical scenarios where true impact parameters are unavailable, evaluation techniques may leverage alternative physics observations or predicted possibility distribution[28, 25]. This work leaves such experimental issues for further investigation in the future.
@@ 3.1 Dataset generation
Datasets are generated following the procedures below: Firstly, collision data is generated, which gives the produced steady particles and their initial kinetic attributes. The physics model of the collision is Isospin-dependent Quantum Molecular Dynamics (IQMD)[35], and the coalescence of nucleons is processed after the IQMD calculation since ion yield is significant in the CEE energy. Then, CeeRoot[36], a platform based on the FairRoot[37] framework, carries out the simulation, which organizes the collision data together with detector geometry and the Geant4 transport model. Finally, we obtain the dataset for model training and testing with proper pre-processing.
During the CeeRoot simulation, the digitization method[38], which translates the Monte-Carlo points of eTOF into the digitized signals, is carefully designed according to the experimental conditions. Processes of detector response, such as noise hits, hit cluster size, time resolution, and time delays, have been considered in digitization. Noise hits are added randomly following the typical MRPC noise level of 1 Hz/cm2. The time of noise hits is obtained by sampling the uniform distribution U(0, 100) ns, where 100 ns is the typical time window for event assembly. The noise hit position is sampled from U(-1.5, 1.5) ns, corresponding to the strip length range. Due to the transverse diffusion of the avalanche charge, hits with multiple fired strips should be considered in the simulation. In digitization, such an effect is simulated by sampling a cluster size distribution measured in detector tests. For each digitized signal, the signal leading time $T_{\rm leading}$ is calculated as follows:
@@
where $t_{\rm hit}$, $x_{\rm hit}$ are the hit time and position, $x_{\rm end}$ is the position of the strip end, $v_{\rm prop}$ is the signal propagation speed along the strip. $\Delta t_{\rm ele}$ is the time delay of the electronics, which is sampled from U(0, 5) ns and constant for each specific channel. $\Delta t_{\rm reso}$ is sampled from a Gaussian distribution with zero mean and 100 ps standard deviation, representing the time resolution. It is inferred from the data normalization that the time delay effect can be eliminated in the model regression and prediction. This is very helpful because of the complexity of the measurement of channel time delay. As the time-slewing effect has not been implemented in the digitization method, the 100 ps smearing is chosen according to the measurement before the time-slewing correction in the cosmic tests[38].
@@
We simulated 1.21 million events of U-U collisions at 0.5 AGeV. The events are distributed uniformly from impact parameter b=0.0 fm to b=12.0 fm with 0.1 fm step. They are randomly split into the train and test set with a 4:1 ratio. Two datasets are extracted from the test set: the so-called $b{\rm d}b$ test set with 120k events and the uniform test set with all the 242k events in the test set. The $b{\rm d}b$ test set is generated following the collision geometry where the differential cross section ${\rm d}\sigma\propto b{\rm d}b$ [^2] Extracted from the test set, the $b{\rm d}b$ test set contains 17 events with $b=0.1$ fm, 33 events with $b=0.2$ fm, …, 2k events with $b=12.0$ fm. It is used to evaluate the overall performances of the models because it is close to the realistic collisions; the uniform test set is used to assess performance as a function of the impact parameter. The train set has 968k events in which 80% are used for model fitting, and 20% are used for early-stop decision (see Section 3.3) after each training epoch.
@@
We collect the features from the total 672 readout strips for each simulated event, including the hit flag, hit time, and hit position. The hit flag is set to 1 if signals are generated from both strip ends. The hit time is calculated as the average of the times recorded at each end, and the hit position corresponds to their half difference. Despite its simplicity, such calculation provides features with explicit physics meaning. The hit flag is set to zero for strips with no hit detected, and the time and position are set to the local average value from the training dataset. The choice of data imputation here is made to obtain zero values after the normalization of the time and position features. The purpose of keeping zero values for the non-fired nodes is to stop the forward propagation of their features implicitly. The first dense layer is designed to exclude the intercept term so that the non-fired nodes with zero input values always result in zero outputs. The absence of a trainable intercept does not affect the capability of the dense layer because the weight on the hit flag plays an equivalent role. Moreover, the zero-value nodes have no function in the attention layer when calculating the linear combination of the nodes with their attributes. Figure 5 shows the counting rate distribution of the 672 readout strips for the eTOF wall. A descending counting rate is shown due to the collision in fix-target mode.
@@
Figure 5: Counting rate distribution of 672 readout strips for eTOF wall.
@@
We generate the phase space dataset with 30$\times$30 binning to compare the model performance with other work. The range of the phase space used for binning is defined as $y_{0}\in$[-0.5,1.5], $p_{\rm T}/m\in$[0,1.5]. Figure 6 shows the cumulative images of the phase space as a function of the impact parameter, where the evolution of the image shape can be easily recognized. However, for a single event, features are less evident due to event-by-event fluctuations, as demonstrated in Figure 6. Two sets of phase space data, with acceptances covered by CEE and only eTOF, have been prepared for a comparison study.
@@
Figure 6: (a) The phase space images in CEE acceptance as a function of the impact parameter. Each image is filled with 10k events with the impact parameter shown in the legend. The binning of the image is 30$\times$30. (b) Image of an event with $b=6$ fm.